OpenStudy (anonymous):

Help :( When hired at a new job selling jewelry, you are given two pay options: Option A: Base salary of $14,000 a year, with a commission of 12% of your sales Option B: Base salary of $19,000 a year, with a commission of 5% of your sales In order for option A to produce a larger income, you would need sell at least $ _____ of jewelry each year. Hint: earning 12% of your sales means you earn $0.12 for every $1.00 sold

OpenStudy (pfenn1):

Let x=sales Option A: Base salary of $14,000 a year, with a commission of 12% of your sales A = 14000 + 0.12 x

OpenStudy (pfenn1):

Can you write the equation for Option B?

OpenStudy (anonymous):

B: Base salary of $19,000 a year, with a commission of 5% of your sales Pay= 19000 + 0.05 (S) ?

OpenStudy (anonymous):

(x)

OpenStudy (pfenn1):

Correct. So now you want to know how much you'd have to sell to make Option A pay more than Option B. \[A \ge B\text{ or }14000+0.12x \ge19000+0.05x\]...

OpenStudy (anonymous):

I got $251,428.571 is that correct /:

OpenStudy (pfenn1):

Wow. That is a lot of jewelry! No. That isn't the amount I calculated. Do you want to show your work?

OpenStudy (anonymous):

1400+0.12x=19000+0.05x -1400 -1400 0.12x=17600 +0.05x -0.05x -0.05x 0.07x=17600 Then i divided both sides by 0.07x what am i doing wrong?

OpenStudy (pfenn1):

1400 should be 14,000

OpenStudy (anonymous):

oh *face palm!

OpenStudy (anonymous):

so is it 71428

OpenStudy (pfenn1):

That agrees with my answer.